OpenStudy (anonymous):

graph the line: -2x+4? help!

OpenStudy (lgbasallote):

y = -2x + 4 i'm going to use the intercept method in graphin when x = 0 y = -2(0) + 4 y = 0 + 4 y = 4 so one of the coordinates would be (0,4) when y = 0 0 = -2x + 4 2x = 4 x = 2 so the other intercept would be (2,0) now plot the points